A small business has an initial investment of
$8176. The unit cost of the product is $11,
and the selling price is $22.
How many units must be sold to break
even?
Answer in units of units

Answer by stanbon(75887)

A small business has an initial investment of
$8176. The unit cost of the product is $11,
and the selling price is $22.
How many units must be sold to break even?
Answer in units
---
Each unit sold yields (22-11) = $11 profit.
---
Break even occurs when x = 8176/11 = 743.27 units are sold.
---
Round up to 744 units.
===========================
Cheers,
Stan H.